OpenStudy (anonymous):

cos75 = (√6+√2)/4 ... correct?

OpenStudy (anonymous):

false

OpenStudy (anonymous):

cos75 = cos45+30 = cos45cos30 - sin45sin30.....

OpenStudy (anonymous):

(sqrt(3)-1)/(2 sqrt(2))

OpenStudy (anonymous):

\[(\sqrt{3}-1)\div(2\sqrt{2})\]

OpenStudy (anonymous):

? i've got cos45 = √ 2/2 , cos30=√ 3/2, sin45 = √ 2/2, sin30=1/2

OpenStudy (anonymous):

diogo has it

OpenStudy (anonymous):

if you rationalize denminator you get \[\frac{\sqrt{6}-\sqrt{2}}{4}\]

OpenStudy (anonymous):

.... confused, √ 2/2 * √ 3/2 = √ 6/4 , and 1/2 * √ 2/2 = √ 2/4, so ...... AH, the symbol is minus, but the numbers are correct, right?

OpenStudy (anonymous):

so my post should read: "cos75 = (√6-√2)/4 .... correct?"

OpenStudy (dumbcow):

yes

OpenStudy (anonymous):

cos75=cos(45+30) =cos45cos30-sin45sin30 =(1/2^.5)(3^.5/2)-(1/2^.5)1/2 (3^.5-1)/2*2^.5 false